# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v1

@@ 3.3 Full Bayesian implementation using MCMC
For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters. Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 3.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform$(0,1)$ priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components. Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP. The local distribution indicators $z_{ij}^{l}$, local distribution proportion $\pi^{l}$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates. However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ stepping-out slice sampling (Neal,, 2003). Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3. For signal detection, we use an FDR-controlled, FNR-optimized hypothesis testing framework based on posterior draws for $\lambda_{ij}$, described next.
@@ 3.4 Signal detection using hypothesis testing
for all AEs $i=1,\dots,I$ and all drugs $j=1,\dots,J$. The null hypothesis $H_{0,ij}$ is rejected in favor of $H_{1,ij}$, identifying the drug-$j$ and AE-$i$ pair to be a signal, if $T_{ij}>k$, where $T_{ij}$ is a test statistic derived from the posterior distribution and $k>1$ is an appropriate threshold. A common choice for $T_{ij}$ is the marginal posterior posterior quantile $T_{ij}(p)$, defined by $\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq T_{ij}(p))=p$ for a small pre-specified quantile level $p\in(0,1)$, where $\Pr{\!}_{\text{posterior}}(\cdot)$ denotes posterior probability given the data, and can be computed via MCMC draws. Hu et al., (2015) suggest using the 5-th posterior quantiles $T_{ij}(0.05)$ as the test statistic with a rejection threshold of $k=2$; however, such deterministic thresholds can be overly conservative (curtailing sensitivity) or overly liberal (failing to control false discoveries). Instead, we consider a family of tests $\{T_{ij}(p)>k\}$ with varying $p$ (reasonably small) and $k>1$, selecting the optimal $(\hat{p},\hat{k})$ that controls the false discovery rate (FDR) while minimizing the false negative rate (FNR). Given $p$ and $k$, FDR and FNR are estimated using the posterior distribution as (Ahmed et al., 2010; Müller et al., 2006; in the expressions below $1(A)$ denotes the indicator of a set $A$, $\hat{=}$ denotes “estimated by”, and $T^{\text{obs}}$ is the computed value of $T$):
@@
where $q_{ij}=\Pr{\!}_{\text{posterior}}(\lambda_{ij}\leq 1)$ is the posterior probability of $H_{0,ij}$, computable from MCMC draws. A grid search is used to determine the optimal $(\hat{p},\hat{k})$ that ensures $\text{FDR}(\hat{p},\hat{k})\leq\alpha$ for a pre-specified level $\alpha$ while minimizing FNR (see Supplement S.3.3 for details on the grids).
@@ 4 SIMULATION ASSESSMENTS
In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings. To generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns. Using a reference $I\times J$ contingency table (described next) from a real-world SRS database, we extract row and column totals $\{\bm{n}_{i\bullet}:i=1,\dots,I\}$ and $\{\bm{n}_{\bullet j}:j=1,\dots,J\}$, respectively, to facilitate random contingency table generation. Next, we generate $\bm{p}_{i\bullet}\sim\text{Dirichlet}(\bm{n}_{i\bullet})$ and $\bm{p}_{\bullet j}\sim\text{Dirichlet}(\bm{n}_{\bullet j})$ for all $i$, $j$ defining the random probability matrix $\bm{p}=((p_{ij}))$ with $p_{ij}\propto\lambda_{ij}^{0}\bm{p}_{i\bullet}\bm{p}_{\bullet j}$ and $\sum_{i,j}p_{ij}=1$. Finally, we generate a random SRS contingency table by sampling from $\text{Multinomial}(\bm{n}_{\bullet\bullet};\bm{p})$, collapsing any generated $n_{ij}=1$ corresponding to a true signal (as defined by $\lambda^{0}_{ij}$) to $n_{ij}=0$ afterward to eliminate signal detection at $n_{ij}=1$.
@@
Reference SRS dataset to compute row and column totals. We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,610 reported cases (Chakraborty et al.,, 2023), to construct our reference SRS contingency table. We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category. Using 46 preferred terms (PTs) for AEs, we form a $47\times 7$ contingency table summarizing the report counts for each drug-AE pair. This dataset has been previously analyzed in Chakraborty et al., (2022).
@@
Constructing the true signal strength matrix $\Lambda^{0}$. To generate $\Lambda^{0}$, we define two signal patterns: (1) fixed row signals, where all drugs (columns) exhibit signals for the same AEs (PTs; rows), and (2) random signals, where signals appear on randomly chosen AEs (different row for each drug). A larger number of fixed row signals increase the between-drug associations, measured via correlations between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ for different statin drugs $j$ and $j^{\prime}$, while more random signals weaken them. The matrix $\Lambda^{0}$ is constructed by specifying fixed row signals and embedding additional random signals in the remaining rows across all statin columns. Signal strengths in these cells range from $1<\lambda_{ij}^{0}\leq 3$ (see Table 1). The remaining cells $(i^{\prime},j^{\prime})$ are designated as non-signals with $\lambda_{i^{\prime}j^{\prime}}^{0}=1$. To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (00 or $0.10$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$. This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables. Within this setup, we designed three simulation scenarios, each with and without zero-inflation (see Table 1).
@@
Table 1: Simulation scenarios describing the construction of the true signal strength matrix $\Lambda^{0}$ encoding signals, non-signals, zero-inflation, and between-drug associations. The first two rows in the table show the number of fixed-AE/row signals and the number of random signals per (statin) column in $\Lambda^{0}$. The zero inflation rate (third row in the table) shows the drug-specific true zero-inflation probability $\omega_{j}^{0}$ in each simulation scenario. The comparator “Other drugs” column has all non-signal $(\lambda_{ij}^{0}=1)$ cells. Finally, the fourth row in the table shows the average Pearson’s correlation coefficient $\rho$ between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ over all statin drug pairs $j$ and $j^{\prime}$ (excluding the “Other drugs” category) as a measure of pairwise dependence between drugs.
@@
No. of fixed row signals 1 1 3 3 6 6
No. of random signals per column 5 5 3 3 1 1
Zero-inflation rate per column 0 0.1 0 0.1 0 0.1
Average $\rho$ between columns 0.04 0.04 0.43 0.43 0.83 0.83
  • The common/fixed PTs (fixed AEs/rows) that are considered signal in each case are as follows:
    • Case 1: Rhabdomyolysis
    • Case 2: Muscle Disorder, Myalgia, Rhabdomyolysis
    • Case 3: Blood Creatine Phosphokinase Increased, Muscle Disorder, Muscle Rupture, Musculoskeletal Discomfort, Myalgia, Rhabdomyolysis
@@
Replicated Data Generation, Model Fitting, and signal detection. For each constructed signal strength matrix $\Lambda$ (Table 1), and using the random contingency table generation mechanism described earlier, we generate $R=1{,}000$ independent random contingency table replicates. In each replicate, we fit our models via MCMC sampling (Supplementary Algorithms S.1, S.2 and S.3) with $10{,}000$ post-burn-in draws after discarding the first $5,000$ burn-in draws; see Supplement S.3 for additional computational details, including starting points. From the computed posteriors, we conduct signal detection using hypothesis testing (Section 3.4).
We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al., (2015) (henceforth DP Hu et al). For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 3.4), with the additional constraint $n_{ij}>1$ for our methods. To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted (Müller et al.,, 2006) posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$. Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT (Chakraborty et al.,, 2022; Huang et al.,, 2011; Ding et al.,, 2020), which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present (see Supplement S.1 for a detailed review of competing approaches).
@@
The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is near zero (Cases 1a,b). Under moderate between-drug association ($\rho=0.43$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$). The gain is even further pronounced under high between-drug association ($\rho=0.83$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$. These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models. In all settings, the FDR and average type I error is strictly controlled for all methods.
The Poisson and ZIP-based DP local-global models perform similarly across all scenarios regardless of the presence of zero-inflated observed counts (sub-cases a and b), due to the flexibility of the DP-based nonparametric framework (see Remark Remark).
All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz. The median computation time per replication for our DP local-global methods (both Poisson and ZIP) was 145 seconds.
@@ 5 REAL DATA ANALYSIS
We use the statin data from Section 4 to demonstrate our method’s utility in real-world SRS data mining. Applying our DP local-global models (using the same MCMC and test statistic generation settings as used for simulated data analysis in Section 4) we compare detected signals against those from the DP Hu et al., BCPNN, GPS, and LRT. Table 2 presents the number of signals identified across statin drugs. The results indicate identical performance for our Poisson and ZIP models, detecting more signals than all competing methods.
@@
Table 2: Summary table of discoveries from our proposed DP local-global models, and comparing them to the other existing methods. The count in each cell represents the number of AE signals detected by the method (along the columns), first separately for each drug and then in total (rows). The number of discoveries made uniquely by our DP Poisson local-global method but not by the other methods are presented in parentheses.
Drug DP Poisson DP ZIP DP Hu et al. BCPNN GPS LRT
Atorvastatin 33 33 (0) 33 (0) 28 (5) 26 (7) 28 (5)
Fluvastatin 14 14 (0) 13 (1) 10 (4) 12 (2) 12 (2)
Lovastatin 12 12 (0) 10 (2) 9 (3) 9 (3) 9 (3)
Pravastatin 23 23 (0) 17 (6) 11 (12) 11 (12) 11 (12)
Rosuvastatin 29 29 (0) 28 (1) 25 (4) 25 (4) 25 (4)
Simvastatin 29 29 (0) 28 (1) 26 (3) 25 (4) 25 (4)
Total 140 140 (0) 129 (11) 110 (31) 110 (32) 110 (30)
To further examine the results, we focus on 20 PTs (AEs/$i$) with the highest average (across statin drugs) posterior medians of $\lambda_{ij}$ from our DP Poisson local-global model (Figure 4). The heatmap indicates non-detection (gray) or the degree of information sharing between drugs, measured via the posterior probabilities of $1-z_{ij}^{l}$ (white-blue gradient; darker blue signifies greater information sharing at each AE). Our method detects several rare AEs with $n_{ij}=2$, some also identified by certain competing methods (though no other method detects all these AEs simultaneously) and some not detected by any competing method (not shown in Figure 4; see Table 2). This, along with strong FDR control, highlights our method’s effectiveness in SRS signal detection.
@@
Figure 4: Heatmap of 20 selected PTs. Cells not detected as signals by our method are colored gray. White-to-blue shades represent the posterior probability of $1-z_{ij}^{l}$, with darker (lighter) blue indicating a higher (lower) probability that the signal strength originates from the global signal strength distribution. The letters “H”, “B”, “G”, and “L” indicate signal detection by the DP Hu et al., BCPNN, GPS, and likelihood ratio test, respectively.
Finally, we focus on inferring structural zeros through the estimated drug-specific zero-inflation probabilities $\omega_{j}$ as returned by our DP ZIP local-global model. For comparison, we also monitor the corresponding estimates provided by the frequentist ZIP (pseudo) LRT method. The two methods yield largely similar results: our approach produces point estimates (posterior mean) for $\omega_{j}$ as 0.02, 0.10, 0.11, 0.08, 0.02, and 0.03 for Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, respectively. The corresponding frequentist maximum likelihood estimates are 0.00, 0.07, 0.16, 0.06, 0.00, and 0.00, respectively.

### v2

@@ 4.3 Full Bayesian implementation using MCMC
For full Bayesian inference on our models, we place independent weakly informative proper priors on all model hyperparameters. Along with the exponential prior for $\alpha_{j}$ and $\alpha$, and half-Cauchy prior for $\beta_{j}$ and $\beta$ as suggested in Section 4.1, we consider a similar Cauchy+(0, $\psi_{\tau}$) prior for $\tau$ with $\psi_{\tau}=1/2$, and independent Uniform(0,1)(0,1) priors for the probability weight parameters, viz., $\pi_{l}$ and $\omega_{j}$ (if included in the model).
We develop efficient MCMC algorithms for rigorous posterior sampling, approximating the DPs using finite mixtures with $K=O(\log(IJ))$ components ishwaran2001gibbs . Each MCMC iteration involves data augmentation to generate latent DP mixture allocations, followed by Gibbs sampling updates for the stick-breaking weights and atoms and the concentration hyperparameters for each local and global DP. The local distribution indicators $z_{ij}^{l}$, local distribution proportion $\pi^{l}$, AE signal strengths $\lambda_{iJ}$ for the reference drug $J$, and zero-inflation indicators $y_{ij}$ and probabilities $\omega_{j}$ (if included in the model) also admit standard Gibbs updates. However, the full conditionals for the DP base Gamma hyperparameters ($\beta_{j}$ and $\beta$) and the reference category shrinkage hyperparameter $\tau$ lack standard forms; for these we employ univariate stepping-out slice sampling neal2003slice steps. Detailed MCMC steps are provided in Supplementary Algorithms S.1, S.2 and S.3. For signal detection, we use an FDR-controlled, FNR-optimized hypothesis testing framework based on posterior draws for $\lambda_{ij}$, described next.
@@ 4.4 Signal detection using hypothesis testing
[…]
@@
[…]
@@ 5 SIMULATION ASSESSMENTS
In this section, we conduct extensive simulations to assess the signal detection performance of our local-global DP Poisson and ZIP models under a wide array of data-generating settings. We follow the SRS contingency table generation approach previously used in the literature chakraborty2022use ; liu2024pattern . Specifically, to generate random SRS contingency tables of dimension $I\times J$, we first define a prespecified true signal strength matrix (described below) $\Lambda^{0}=((\lambda_{ij}^{0}))$ encoding signals, structural zeros, non-signals, and between-drug dependency patterns. Using a reference $I\times J$ contingency table (described next) from a real-world SRS database, we extract row and column totals $\{\bm{n}_{i\bullet}:i=1,\dots,I\}$ and $\{\bm{n}_{\bullet j}:j=1,\dots,J\}$, respectively, to facilitate random contingency table generation. Next, we generate $\bm{p}_{i\bullet}\sim\text{Dirichlet}(\bm{n}_{i\bullet})$ and $\bm{p}_{\bullet j}\sim\text{Dirichlet}(\bm{n}_{\bullet j})$ for all $i$, $j$ defining the random probability matrix $\bm{p}=((p_{ij}))$ with $p_{ij}\propto\lambda_{ij}^{0}\ \bm{p}_{i\bullet}\ \bm{p}_{\bullet j}$ and $\sum_{i,j}p_{ij}=1$. Finally, we generate a random SRS contingency table by sampling from $\text{Multinomial}(\bm{n}_{\bullet\bullet};\bm{p})$, collapsing any generated $n_{ij}=1$ corresponding to a true signal (as defined by $\lambda^{0}_{ij}$) to $n_{ij}=0$ afterward to eliminate signal detection at $n_{ij}=1$.
Reference SRS dataset to compute row and column totals
@@
We use the FDA FAERS database (Q3 2014-Q4 2020), which catalogs a total of 63,976,61063{,}976{,}610 reported cases chakraborty2023likelihood , to construct our reference SRS contingency table. We focus on six statin medications: Atorvastatin, Fluvastatin, Lovastatin, Pravastatin, Rosuvastatin, and Simvastatin, collapsing all other drugs into a comparator “Other drugs” category. For expository purposes, and to create a sufficiently large SRS contingency table, we focused on 1491 preferred terms (PTs) as AEs that are common (appear at least once) for these statin drugs as well for a separate group of drugs unrelated to statin, namely, Gadolinium-based contrast agents (GBCA) drugs chakraborty2022use . All other PTs recorded in the database within this period were collapsed into a heterogeneous category labeled “Other AE”, yielding a $1492\times 7$ contingency table summarizing the report counts for each drug-AE pair.
Constructing the true signal strength matrix $\Lambda^{0}$
@@
To generate $\Lambda^{0}$, we define two signal patterns: (1) fixed row signals, where all drugs (columns) exhibit signals for the same AEs (PTs; rows), and (2) random signals, where signals appear on randomly chosen AEs (different row for each drug). A larger number of fixed row signals increase the between-drug associations, measured via correlations between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ for different statin drugs $j$ and $j^{\prime}$, while more random signals weaken them. The matrix $\Lambda^{0}$ is constructed by specifying fixed row signals and embedding additional random signals in the remaining rows across all statin columns. Signal strengths in these cells range from $1<\lambda_{ij}^{0}\leq 3$ (see Table 1). The remaining cells $(i^{\prime},j^{\prime})$ are designated as non-signals with $\lambda_{i^{\prime}j^{\prime}}^{0}=1$. To introduce zero inflation, we prespecify true $\omega_{j}^{0}$ (0 or $0.30$; Table 1), and set $\omega_{j}^{0}I$ randomly selected non-signal cells to $\lambda_{ij}^{0}=0$. This approach yields a controlled $\Lambda^{0}$ to generate random SRS contingency tables. Within this setup, we designed three simulation scenarios, each with and without zero-inflation (see Table 1).
@@
Table 1: Simulation scenarios describing the construction of the true signal strength matrix $\Lambda^{0}$ encoding signals, non-signals, zero-inflation, and between-drug associations. The first two rows in the table show the number of fixed-AE/row signals and the number of random signals per (statin) column in $\Lambda^{0}$. The zero inflation rate (third row in the table) shows the drug-specific true zero-inflation probability $\omega_{j}^{0}$ in each simulation scenario. The comparator “Other drugs” column has all non-signal $(\lambda_{ij}^{0}=1)$ cells. Finally, the fourth row in the table shows the average Kendall’s $\tau$ between $\lambda_{ij}^{0}$ and $\lambda_{ij^{\prime}}^{0}$ across all statin drug pairs $j$ and $j^{\prime}$ (excluding the “Other drugs” category) as a measure of pairwise association of their AE signal strengths.
@@
No. of fixed row signals 5 5 10 10 20 20
No. of random signals per column 17 17 10 10 3 3
Zero-inflation rate per column 0 0.3 0 0.3 0 0.3
Average Kendall’s $\tau$ between columns 0.12 0.12 0.49 0.49 0.77 0.77
@@
Replicated Data Generation, Model Fitting, and signal detection
For each constructed signal strength matrix $\Lambda$ (Table 1), and using the random contingency table generation mechanism described earlier, we generate R=1, [^0] R=1{,}000 independent random contingency table replicates. In each replicate, we fit our models via MCMC sampling (Supplementary Algorithms S.1, S.2 and S.3) with 10,00010{,}000 post-burn-in draws after discarding the first $5,000$ burn-in draws; see Supplement S.2 for additional computational details, including starting points for the MCMC algorithms. From the computed posteriors, we conduct signal detection using hypothesis testing (Section 4.4).
We compare our signal detection performance against existing Bayesian SRS data mining methods, viz., BCPNN, GPS, and the local-only DP model of Hu et al.hu2015signal (henceforth DP Hu et al). For our models and DP Hu et al., signal detection uses an optimized posterior quantile level $\hat{p}$ (fixed at 0.05 for DP Hu et al.) and threshold $\hat{k}$ ensuring FDR $\leq 0.05$ while minimizing FNR (Section 4.4), with the additional constraint $n_{ij}>1$ for our methods. To ensure fair FDR control ($\leq 0.05$) in BCPNN and GPS, we leverage the independence of signal strengths in the (empirical) Bayes posteriors for these two models and compute Benjamini-Hochberg-adjusted muller2006fdr posterior probabilities of being a non-signal; a drug-$j$ and AE-$i$ pair is then classified as a signal if the adjusted probability is $\leq 0.05$. Additionally, we evaluate the state-of-the-art frequentist ZIP-based extended (pseudo) LRT chakraborty2022use ; huang2011likelihood ; ding2020evaluation , which ensures FDR control ($\leq 0.05$) with high sensitivity while addressing zero inflation in the counts when present.
@@
The first two rows of Figure 3 show comparable (when $\lambda^{0,\text{signal}}<1.5$) or slightly better (when $\lambda^{0,\text{signal}}\geq 1.5$) sensitivity for our local-global methods compared to all other methods when the between-drug association is very small ($\tau=0.12$; Cases 1a,b). Under moderate between-drug association ($\tau=0.49$; Cases 2a,b), our methods perform noticeably better in detecting more true signals, achieving up to twice the sensitivity, particularly for weak or moderate true signal strengths ($1.1\leq\lambda^{0,\text{signal}}\leq 1.7$). The gain is even further pronounced under high between-drug association ($\tau=0.77$; Cases 3a,b), where our methods attain sensitivities of $\geq 0.70$ with only moderate true signal strength ($\lambda^{0,\text{signal}}\geq 1.7$), while all competitors fail to reach $0.70$ sensitivity even for $\lambda^{0,\text{signal}}=3$. These results highlight the substantial boost in signal detection from effectively leveraging between-drug similarities as facilitated by our models. In all settings, the FDR and average type I error is strictly controlled for all methods.
The Poisson and ZIP-based DP local-global models perform similarly across all scenarios regardless of the presence of zero-inflated observed counts (sub-cases a and b), due to the flexibility of the DP-based nonparametric framework (see Section 4.1).
All computations were performed on the cluster computing facility at the Center for Computational Research at the University at Buffalo, with each process running on 16 GB RAM and a clock speed of 1.6-3 GHz. The median computation time per replication for our DP local-global methods (both Poisson and ZIP) is 48 minutes.
@@ 6 REAL DATA ANALYSIS
We use the statin data from Section 5 to demonstrate our method’s utility in real-world SRS data mining. We fit our DP-Poisson and DP-ZIP local-global models to the dataset, setting the finite mixture approximation of the DP with $K=20$ components, and using the same MCMC and test statistic generation settings as used for simulated data analysis in Section 5). We compare the detected signals against those from the DP Hu et al., BCPNN, GPS, and LRT. Table 2 presents the number of signals identified across statin drugs by all these methods. The results indicate very similar performance for our Poisson and ZIP local-global models, both of which detect a substantially larger number of signals than all competing methods.
@@
Table 2: Summary table of discoveries from our proposed DP local-global models, and comparing them to the other existing methods. The count in each cell represents the number of AE signals detected by the method (along the columns), first separately for each drug and then in total (rows). The numbers in parentheses represent the differences in the number of discoveries made by each method when compared to our DP Poisson local-global method.
Drug DP Poisson local-global DP ZIP local-global DP Hu et al. BCPNN GPS LRT
Atorvastatin 308 306 (2) 307 (1) 234 (74) 203 (105) 126 (182)
Fluvastatin 136 136 (0) 122 (14) 68 (68) 60 (76) 26 (110)
Lovastatin 90 94 (-4) 59 (31) 37 (53) 30 (60) 11 (79)
Pravastatin 226 227 (-1) 210 (16) 122 (104) 91 (135) 35 (191)
Rosuvastatin 314 314 (0) 297 (17) 228 (86) 195 (119) 110 (204)
Simvastatin 314 317 (-3) 257 (57) 220 (94) 183 (131) 94 (220)
Total 1388 1394 (-6) 1252 (136) 909 (479) 762 (626) 402 (986)
To further examine the signal detection results, we focus on our DP Poisson local-global model and zoom in on 15 selected PTs, whose signal strength estimates are plotted using heatmaps in Figure 4. The figure is divided into three panels based on the frequencies of the AEs, and the colors in the heatmaps indicate either non-detection (gray) or the estimated signal strengths measured through the posterior medians of $\lambda_{ij}$ (white-to-blue gradient with a darker blue signifying a stronger signal strength). Panel A displays the five most common AEs with the largest frequencies (total counts for the AEs across all six statin drugs being greater than $2000$) and the strongest estimated signal strengths (having the largest posterior medians of $\lambda_{ij}$ across all statin drugs $j$) that are detected as a signal for at least one statin drug by our method. The AEs displayed in this panel are all well-documented muscle-related conditions, such as rhabdomyolysis antons2006clinical ; tournadre2020statins and myopathy ballantyne2003risk ; abd2011statin , and their detection by our method is consistent with the performance of all competing methods. Panel B contains five semi-rare AEs (with observed counts across statin drugs ranging between 500 and 2000) that exhibit the strongest estimated signal strengths. This category primarily includes some specific types of disorders, such as nervous system disorders, and interestingly, several of these detected signals are not identified by any competing method for the newer or less frequently used statin drugs. For example, unlike other methods, our model detects lovastatin-induced drug intolerance as a signal, which has been flagged as a potential side effect of lovastatinahmad2014statin . Additionally, our method detects signals for paraesthesia and hypoaesthesia, symptoms of peripheral neuropathy, associated with lovastatin use; this finding is also consistent with previous research chong2004statin ; wang2021association . By appropriately borrowing strengths from the more commonly used statins, our method is successfully able to identify potential lovastatin-induced AE signals in these specific examples. Finally, Panel C focuses on five rare AEs (with observed counts across statin drugs being lower than 500) with the strongest estimated signal strengths. Within this group, our method identifies several AEs related to neurological and movement issues, such as movement disorder and swollen tongue, and some other mixed disorders. Many of these AEs were not originally discovered in clinical trials, but later were flagged as potential signals in specific, targeted pharmacovigilance investigations. For example, fluvastatin-induced swollen tongue, also known as angioedema, has been recently recorded kumar2024fluvastatin . Similar to the semi-rare AEs displayed in Panel B, our method is able to identify these rare AE signals for the less frequently used drug fluvastatin by appropriately sharing AE signal strength information across all statin drugs.
@@
Figure 4: Heatmaps of 15 selected PTs. Cells not detected as signals by our method are colored gray. White-to-blue shades represent the median of the posterior of signal strength parameter $\lambda_{ij}$, with darker (lighter) blue indicating a stronger (weaker) signal strength. Panel A contains 5 common PTs among all statin drugs. Panel B contains 5 semi-rare PTs with new discovered signals by our proposed method. Panel C contains 5 rare PTs with new discovered signals by our proposed method. Count $n$, expected count $E$, and the 90% credible interval of the posterior distribution of $\lambda_{ij}$ are shown in each cell. The letters “H”, “B”, “G”, and “L” indicate signal detection by the DP Hu et al., BCPNN, GPS, and likelihood ratio test, respectively.